# Improving Our Knowledge of the Solar Near-Surface Shear Layer: The Special Case of the Leptocline

### v1

@@ 1 Introduction
In recent years, numerous studies have focused on the physical conditions prevailing in the Sun’s subsurface layers for at least two reasons. The first addresses the problem of how both the physical conditions in subsurface layers of the Sun and the nature of the magnetic flux tubes of active regions are reflected in the structure and behavior of these regions at the surface (e.g. Howard 1996; Choudhuri and Jha 2023; Rabello Soares et al. 2024; Kitchatinov 2023; Vasil et al. 2024). The second relates to differential rotation: the aim is to understand how the solar rotation, which is not uniform in latitude, also varies in depth and time. One of the most intriguing features is the rotation rate, which is faster at the equator than at the pole, a phenomenon known as “differential rotation” (for instance and references therein Javaraiah and Gokhale 2002; Javaraiah 2003; Tassoul 2000; Howe 2020).
@@
Helioseismic studies, which are a powerful tool for probing the solar interior in three dimensions, show significant velocity variations in the near-surface layers. The treatment of the superadiabiatic region supposes a proper description of turbulent convection and detailed radiative energy transport and thermodynamic calculations. We also need to understand how the turbulent convection interacts with solar rotation. Furthermore, helioseismic studies have illustrated that the most significant changes with the solar cycle occur in a near-surface shear layer (NSSL), occupying around 5% of the solar radius at the top of the convection zone. The velocity shear may convert a part of the poloidal magnetic field into the toroidal field, and, in addition to the global dynamo operating in the bulk of the convection zone (e.g. Pipin et al. 2023), the magneto-rotational instability may play a certain role (Vasil et al. 2024).
@@ 2 Radial Gradients of Solar Rotation: Tachocline and Leptocline
The internal rotation of the Sun has been observed almost uninterruptedly since 1996 from two space missions, the Solar and Heliospheric Observatory (SoHO) (Scherrer et al. 1995) and the Solar Dynamics Observatory (SDO) (Scherrer et al. 2012) as well as from the ground-based Global Oscillation Network Group (GONG) network (Hill et al. 1996). The rotation rate is inferred by applying helioseismic inversion techniques to the rotational splitting of solar oscillation frequencies measured every 72 days from Doppler velocity images (Schou et al. 1997).
@@
These results are generally consistent with previous inferences of the internal solar rotation (e.g. Thompson et al. 1996; Kosovichev et al. 1997; Schou et al. 1998). In particular, the results reveal two zones of sharp gradients at the bottom of the convection zone (the tachocline) and its top (the Near-Surface Shear Layer, NSSL). The gradient maximum values at the equator and 60 degrees latitude, connected by a dashed line in Fig. 1d, indicated that the tachocline is deeper at the equator than at the high latitude by about 0.02-0.03 $R_{\odot}$. This means the tachocline has a prolate shape as initially argued by Gough and Kosovichev (1995).
@@
  • In a shallow layer, between 0.96 $R_{\odot}$ and 1 $R_{\odot}$, the rotation rate decreases by about 5% at all latitudes, showing however a more complex behavior near the surface. This layer is called the Near Subsurface Shear Layer (NSSL).
@@
Figure 2: The radial gradient of the mean rotation rate, averaged over all SoHO/MDI and SDO/HMI measurements in 1996-2024, as a function of depth in the near-surface shear layer at 0, 30, and 60 latitudes.
@@
Figure 3: Variations of the rotational gradient, $d\log\Omega/d\log r$, in the leptocline a) with the depth below the solar surface at three latitudes, and b) with latitude at three depths (shown in the figure); c) latitude-radius diagram of the rotational gradient in the NSSL and leptocline.
@@ 3 Solar-Cycle Variations of the Rotational Gradient
Figure 4: a) The magnetic time-latitude (“butterfly”) diagram for the radial component of the line-of-sight magnetic field from the SoHO and SDO data; b-c) the time-latitude diagrams for the rotational gradient, $d\log\Omega/d\log r$ below the leptocline ($r/R=0.97-0.99$) and in the leptocline ($r/R=0.99-1.00$) respectively.
@@
To illustrate the solar-cycle variations in the NSSL and leptocline, in Figure 4 we present the variations of the radial gradient as a function of time and latitude in two layers, just below the leptocline, at $r/R_{\odot}=0.97-0.99$ (or in the depth range of 7-21 Mm) and in the leptocline (panel b), at $r/R_{\odot}=0.99-1.00$ (the corresponding depth range is 0-7 Mm (panel c). The comparison with the corresponding magnetic butterfly diagram (panel a) shows that below the leptocline, the negative gradient becomes stronger (dark blue areas in panel b) in the strong magnetic field regions migrating toward the equator and weaker in the high-latitude regions during the sunspot cycles. In the leptocline (panel c), the variations are less pronounced and have more complicated structures, which, however, resemble the overlapping extended cycles of the torsional oscillations. In particular, the rotational gradient is stronger not only during the activity maxima but also during the activity minima when there are no strong magnetic fields on the solar surface. The gradient enhancement in the leptocline in quiet-Sun regions was previously noticed in the ring-diagram data (Komm 2022).
@@ 4 Variations of the Helioseismic Radius of the Sun With Respect to the Leptocline
Figure 5: Variations of the seismic radius of subsurface layers during Solar Cycles 23 and 24: a) the sunspot number of these cycles; b) the time-depth diagram of subsurface displacements, $\delta r$, inferred from f-mode frequencies obtained from the SoHO and SDO data; c) the variations of the Sun’s seismic radius obtained by averaging the displacements of the depth; d) the variations of the seismic radius ($\delta r$) in the leptocline, at a depth of 5 Mm. The data are adapted from Kosovichev and Rozelot (2018).
Further observations revealed variations of the seismic radius with the solar cycle, resulting in a reduction by several km with an increase in solar activity (Dziembowski et al. 1998). Using nine years of data from SoHO, Lefebvre and Kosovichev (2005) established a variability of the helioseismic radius in antiphase with the solar activity, decreasing by about 2 km at the solar maximum.
@@ 5 Radiative Hydrodynamics Simulations of the Leptocline
Kitiashvili et al. (2023) analyzed realistic 3D radiative hydrodynamics simulations of solar subsurface dynamics in the presence of rotation in a local domain 80 Mm wide and 25 Mm deep, located at 30 degrees latitude. The simulations revealed the development of a shallow 8-Mm deep substructure of the Near-Surface Shear Layer (NSSL), characterized by strong turbulent flows and radial rotational gradient corresponding to the leptocline (Fig. 6). It is located in the hydrogen ionization zone associated with enhanced anisotropic overshooting of convective flows (revealed by enhanced fluctuations of density, $\rho^{\prime}_{\rm RMS}$ in Fig. 6b) into a less convectively unstable layer at a depth of about 8-12 Mm between the H I/He I and He II ionization zones, as illustrated by the adiabatic exponent $\Gamma_{1}$ in Fig. 6a.
@@ 6 Conclusions
In summary, the results of global and local helioseismology as well as 3D radiative hydrodynamic simulations show that the near-surface rotational shear layer (NSSL) occupying the top 15% of the solar convection zone, the depth range $\approx 30-35$ Mm) has a distinct substructure, the leptocline, which is about 8 Mm deep and characterized by enhanced turbulent convection and a sharp increase in the rotational shear.

### v2

@@ Abstract
V1: Figures and results unchanged. Minor changes: corrected typos; added references; acronym NSSL used instead full denomination.
@@ 1 Introduction
In recent years, numerous studies have focused on the physical conditions prevailing in the Sun’s subsurface layers for at least two reasons. The first addresses the problem of how both the physical conditions in subsurface layers of the Sun and the nature of the magnetic flux tubes of active regions are reflected in the structure and behavior of these regions at the surface (e.g. Howard 1996; Choudhuri and Jha 2023; Rabello Soares et al. 2024; Kitchatinov 2023; Vasil et al. 2024). The second relates to differential rotation: the aim is to understand how the solar rotation, which is not uniform in latitude being faster at the equator than at the pole, varies in depth and time, a phenomenon known as “differential rotation” (for instance and references therein Javaraiah and Gokhale 2002; Javaraiah 2003; Tassoul 2000; Howe 2020).
@@
Helioseismic studies, which are a powerful tool for probing the solar interior in three dimensions, show significant velocity variations in the near-surface layers (see for instance Komm 2022). Density decreases monotonically by several orders of magnitude; other physical conditions play an important role (see for Lefebvre et al. 2009). The treatment of the superadiabiatic region supposes a proper description of turbulent convection and detailed radiative energy transport and thermodynamic calculations, and we need to understand how the turbulent convection interacts with solar rotation.
Furthermore, helioseismic studies have illustrated that the most significant changes with the solar cycle occur in a Near-Surface Shear Layer -hereinafter called NSSL,occupying around 5% of the solar radius at the top of the convection zone. The velocity shear may convert a part of the poloidal magnetic field into the toroidal field, and, in addition to the global dynamo operating in the bulk of the convection zone (e.g. Pipin et al. 2023), the magneto-rotational instability may play a certain role (Vasil et al. 2024).
@@ 2 Radial Gradients of Solar Rotation: Tachocline and Leptocline
The internal rotation of the Sun has been observed almost uninterruptedly since 1996 from two space missions, the Solar and Heliospheric Observatory (SoHO; Scherrer et al. 1995) and the Solar Dynamics Observatory (SDO; Scherrer et al. 2012) as well as from the ground-based Global Oscillation Network Group (GONG) network (Hill et al. 1996). The rotation rate is inferred by applying helioseismic inversion techniques to the rotational splitting of solar oscillation frequencies measured every 72 days from Doppler velocity images (Schou et al. 1997).
@@
These results are generally consistent with previous inferences of the internal solar rotation (e.g. Thompson et al. 1996; Kosovichev et al. 1997; Schou et al. 1998). In particular, the results reveal two zones of sharp gradients at the bottom of the convection zone (the tachocline) and its top (the NSSL). The gradient maximum values at the equator and 60 degrees latitude, connected by a dashed line in Fig. 1d, indicated that the tachocline is deeper at the equator than at the high latitude by about 0.02-0.03 $R_{\odot}$. This means the tachocline has a prolate shape as initially argued by Gough and Kosovichev (1995).
@@
  • In a shallow layer, between 0.96 $R_{\odot}$ and 1 $R_{\odot}$, the rotation rate decreases by about 5% at all latitudes, showing however a more complex behavior near the surface. This layer is called the NSSL.
@@
Figure 2: The radial gradient of the mean rotation rate, averaged over all SoHO/MDI and SDO/HMI measurements in 1996-2024, as a function of depth in the Near-Surface Shear Layer (NSSL) at three latitudes indicated in the figure.
@@
Figure 3: Variations of the rotational gradient, $d\log\Omega/d\log r$, in the leptocline a) with the depth below the solar surface at three latitudes, b) with latitude at three depths (shown in the figure); c) with latitude-radius diagram of the rotational gradient in the NSSL (0.96-1.0 $R_{\odot}$) and leptocline (0.99-1.0 $R_{\odot}$).
@@ 3 Solar-Cycle Variations of the Rotational Gradient
Figure 4: a) The magnetic butterfly diagram for the radial component of magnetic field calculated from the SoHO and SDO line-of-sight magnetic field data, assuming that the magnetic field on the solar surface is predominately radial; b-c) the time-latitude diagrams for the rotational gradient, $d\log\Omega/d\log r$ below the leptocline ($r/R_{\odot}=0.97-0.99$) and in the leptocline ($r/R_{\odot}=0.99-1.00$) respectively.
@@
To illustrate the solar-cycle variations in the NSSL and leptocline, in Figure 4 we present the variations of the radial gradient as a function of time and latitude in two layers, just below the leptocline, at $r/R_{\odot}=0.97-0.99$ (or in the depth range of 7-21 Mm) and in the leptocline (panel b), at $r/R_{\odot}=0.99-1.00$ (the corresponding depth range is 0-7 Mm; panel c). The comparison with the corresponding magnetic butterfly diagram (panel a) shows that below the leptocline, the negative gradient becomes stronger (dark blue areas in panel b) in the strong magnetic field regions migrating toward the equator and weaker in the high-latitude regions during the sunspot cycles. In the leptocline (panel c), the variations are less pronounced and have a more complicated structure, which, however, resembles the overlapping variation of the extended cycle of the torsional oscillations. In particular, the rotational gradient is stronger not only during the activity maxima but also during the activity minima when there are no strong magnetic fields on the solar surface. The gradient enhancement in the leptocline in quiet-Sun regions was previously noticed in the ring-diagram data (Komm 2022).
@@ 4 Variations of the Helioseismic Radius of the Sun With Respect to the Leptocline
Figure 5: Variations of the seismic radius of subsurface layers during Solar Cycles 23 and 24: a) the sunspot number of these cycles; b) the time-depth diagram of subsurface displacements, $\delta r$, inferred from f-mode frequencies obtained from the SoHO and SDO data; c) the variations of the Sun’s seismic radius obtained by averaging the displacements of the depth; d) the variations of the seismic radius ($\delta r$) in the leptocline, at a depth of 5 Mm. The data are adapted from Kosovichev and Rozelot (2018). Note that the averaged seismic radius changes in anti-correlation, meaning that it becomes smaller during the solar maxima, decreasing by about 3 km during the Cycle 23 maximum in 2000-2003, and by about 2 km during the Cycle 24 maximum in 2013-2015.
Further observations revealed variations of the seismic radius with the solar cycle, resulting in a reduction by several km with an increase in solar activity (Dziembowski et al. 1998). Using nine years of data from SoHO, Lefebvre and Kosovichev (2005) established a variability of the helioseismic radius in antiphase with the solar activity, decreasing by about 2 km at the solar maximum (Figure 5). Indeed, observations of the solar radiation integrated over the entire solar spectrum (total irradiance), obtained by space-based experiments over the last decades, have demonstrated that total irradiance varies on time scales of minutes to the 11-year solar cycle (or more). As pointed out by Pap et al. (2001), if the central energy source remains constant while the rate of energy emission from the surface varies, there must be an intermediate reservoir where the energy can be stored or released depending on the variable rate of energy transport. The solar gravitational and magnetic fields could act as such a mechanism. If the energy is stored in this energy reservoir, it will result in a change in the solar radius, rending such faint variations plausible, by considering that the energy can be stored or released, depending on the variable rate of energy transport (Pap et al. 1998).
@@ 5 Radiative Hydrodynamics Simulations of the Leptocline
Kitiashvili et al. (2023) analyzed realistic 3D radiative hydrodynamics simulations of solar subsurface dynamics in the presence of rotation in a local domain 80 Mm wide and 25 Mm deep, located at 30 degrees latitude. The simulations revealed the development of a shallow 8-Mm deep substructure of the NSSL, characterized by strong turbulent flows and radial rotational gradient corresponding to the leptocline (Fig. 6). It is located in the hydrogen ionization zone associated with enhanced anisotropic overshooting of convective flows (revealed by enhanced fluctuations of density, $\rho^{\prime}_{\rm RMS}$ in Fig. 6b) into a less convective unstable layer at a depth of about 8-12 Mm between the H I/He I and He II ionization zones, as illustrated by the adiabatic exponent $\Gamma_{1}$ in Fig. 6a.
@@ 6 Conclusions
In summary, the results of global and local helioseismology as well as 3D radiative hydrodynamic simulations show that the NSSL (Near Sub Surface Layer) occupying the top 15% of the solar convection zone, the depth range $\approx 30-35$ Mm) has a distinct substructure, the leptocline, which is about 8 Mm deep and characterized by enhanced turbulent convection and a sharp increase in the rotational shear.